OpenStudy (anonymous):

Write the equation of a line that is perpendicular to the given line and that passes through the given point. 2x + 4y = –6; (2, 5) y = -2x + 1 y = 1/2x – 8 y = 2x + 1 y = 1/2x + 1

OpenStudy (mathmale):

Question Number One: how do we find the slope of this given line? Question Number Two: how do we find the slope of a line perpendicular to the given one? Question Number Three: What's the point-slope formula for a straight line, and why would it be preferable to use that equation?

OpenStudy (anonymous):

2 y=mx+b? y - y1 = m(x - x1)

OpenStudy (anonymous):

It is the 3rd one?